I would like an algebra expression for the following problem:
A number increased by 8 is the difference of the number and a product of the number multiplied by-4.

Answer by WannabeCAgirl83(35)

the unknown number: x
the unknown number increased by 8: x + 8
a product of the unknown number multiplied by -4: - 4x

If x + 8 is the difference between x and - 4x, then x + (x + 8) is equal to - 4x, so...

x + (x + 8) = - 4x | re-arranging
2x + 8 = - 4x | - 2x on both sides
8 = -6x | : -6 on both sides
- 1 1/3 = x